# Mixed-wet percolation on a dual square lattice

### v1

@@ 4.4 Fractal dimension
Figure 11: Plot of mass of the spanning perimeter bond cluster $B_{\rm max}$ (on the dual lattice) against $L$ in red circles and plot of mass of the nnn spanning site cluster $S_{\rm max}$ (on the primal lattice) against $L$ in blue circles. The red and the blue lines of slope $91/48$ are guides to the eye.
Since the exponents $\tau$ and $\sigma$ of the cluster size distribution are the same as that of the ordinary site percolation, all other exponents related to different moments of the $n_{b}(p)$ would be the same. As a result, the fractal dimension is expected to be the same as that of the ordinary site percolation. However, we present here an independent measurement of the fractal dimension of the spanning perimeter bond clusters. A typical spanning perimeter bond cluster (in red) is shown in Fig.4(a). The spanning cluster looks very different from a spanning ordinary bond percolation cluster in Fig.4(b). However, the cluster has holes of all possible sizes. It is tortuous and looks self-similar. The mass of the spanning perimeter bond clusters at $p=p_{c}$ is supposed to scale with the system size $L$ as $B_{\rm max}(p_{c},L)\sim L^{d_{f}}$. Note that $B_{\rm max}(\rho_{c},L)$ and $B_{\rm max}(p_{c},L)$ are the same as they correspond to the same critical point. In Fig. 11, the average mass $B_{\rm max}$ of the spanning cluster (shown in black circles) is plotted against $L$ in the log-log scale. The black straight line of slope $d_{f}=91/48$ is a guide to the eye. The data points follow the black straight line quite nicely. The blue squares represent the average mass $S_{\rm max}$ of the spanning clusters of occupied sites on the primal lattice. Those site clusters represent the nnn site percolation spanning clusters. The blue line is parallel to the red line and has a slope $d_{f}=91/48$. Thus, the fractal dimension $d_{f}$ of the mixed wet percolation spanning cluster and that of the nnn site percolation are the same as that of the ordinary (nn) site percolation spanning cluster. It can be seen that $B_{\rm max}(L)>S_{\rm max}(L)$ as $\rho_{c}>p_{c}$. Therefore, even if there is a mass difference at a given $L$, both of them scale with the same fractal dimension. Since $d_{f}$ and other exponents are the same as that of percolation, all the hyper-scaling relations are satisfied for the perimeter bond cluster percolation.
@@ 4.5 Random bond mixed wet percolation (RBMWP) model and results
In Fig. 12, we present a typical morphology of RBMWP on $100\times 100$ square dual lattice at $p=0.5$ and $\omega=0.935$. The RBMWP clusters do not look very different from the mixed wet percolation clusters shown in Fig. 4 except for some dangling bonds. The clusters in RBMWP need not always be a combination of closed loops.
Refer to caption Refer to caption
(a) (b)
Figure 13: (a) Plot of $\omega_{c}(L)$ versus $L^{-1/\nu^{\prime}}$ taking $\nu^{\prime}=4/3$, the correlation length exponent of percolation. This is done for values of $p=0.44,0.47$ and $0.50$. (b) Plot of $\omega_{c}$ for various values of $p$ between $p_{c}$ and $1-p_{c}$.
@@
$\Pi_{b}(\omega,L)=\widetilde{\Pi}_{b}\left[(\omega-\omega_{c})L^{1/\nu^{\prime}}\right]$ (24)
@@
$\Pi_{b}(\omega_{c}(L),L)=\widetilde{\Pi}_{b}\left[(\omega_{c}(L)-\omega_{c})L^{1/\nu^{\prime}}\right]=1/2$ (25)
@@
$\omega_{c}(L)=\omega_{c}+AL^{-1/\nu^{\prime}}$ (26)
@@
Figure 14: Plot of $\Pi^{\prime}_{b}(\omega_{c},L)$ versus $L$ in double logarithmic scale for $p=0.44,0.47$ and $0.50$. The correlation length exponent is found as $1/\nu^{\prime}\approx 0.74\pm 0.004$ for all the three cases.
@@
$\Pi^{\prime}_{b}(\omega_{c},L)=aL^{1/\nu^{\prime}}$ (27)
where $a=\widetilde{\Pi}^{\prime}_{b}[0]$ and the prime () indicates the derivative with respect to $\omega$. The derivatives are measured at the critical thresholds $\omega_{c}$ by fitting a third-degree polynomial around $\omega_{c}$ for each $L$. In Fig. 14, $\Pi^{\prime}_{b}$ are plotted against $L$ in a double logarithmic scale. The slope that indicates $1/\nu^{\prime}$ was determined by linear regression and was found to be $1/\nu^{\prime}=0.740\pm 0.004$. That is very close to that of the ordinary percolation one, i.e.; $\nu^{\prime}=\nu=4/3$.
Refer to caption Refer to caption
(a) (b)
Figure 15: (a) Plot of cluster size distribution $n_{b}(\omega_{c})$ against the size of the clusters $b$ at the threshold $\omega_{c}=0.9350$ for $p=0.5$ on a double logarithmic scale for a lattice size $L=1000$. The red line of slope $91/48$ is a guide to the eye. (b) Plot of the scaled cluster size distribution $n_{b}(\omega)/n_{b}(\omega_{c})$ against the scaled variable $(\omega-\omega_{c})b^{\sigma}$ for $p=0.5$ on a lattice of size $L=2000$. A good collapse of data is obtained for $\sigma^{\prime}=\sigma=36/91$.
@@
$n_{b}(\omega)=b^{-\tau^{\prime}}\mathcal{N}_{b}\left[(\omega-\omega_{c})b^{\sigma^{\prime}}\right]$ (28)
@@
$n_{b}(\omega_{c})\sim b^{-\tau^{\prime}}$ (29)
where $b$ is the size of the clusters and $\tau^{\prime}$ is the cluster size distribution exponent. In Fig. 15 (a), $n_{b}(\omega_{c})$, for $p=0.5$ and $\omega_{c}=0.9350$, is plotted against the cluster size $b$ on a double logarithmic scale. The exponent $\tau^{\prime}$ measured by linear regression of the data points is found to be $\tau^{\prime}=2.058\pm 0.005$. It is very close to the exponent $\tau=187/91$ as that of ordinary site percolation, shown by a red line as a guide to the eye. It seems $\tau^{\prime}$ of RBMWP is the same as that $\tau$ of ordinary site percolation as well as mixed wet percolation.
@@
$\frac{n_{b}(\omega)}{n_{b}(\omega_{c})}\sim\mathcal{N}_{b}\left[(\omega-\omega_{c})b^{\sigma^{\prime}}\right]$ (30)
@@
Figure 16: Plot of mass of the spanning cluster $M_{b,\rm max}$ against lattice size $L$ on a double logarithmic scale. The black straight line is the line of linear regression with slope $d_{f}=1.895\pm 0.004$.
Although $\nu^{\prime}$ and $\tau^{\prime}$ were very close to corresponding exponents of ordinary site percolation, we will verify the fractal dimension of the spanning clusters of RBMWP. A typical spanning cluster of RBMWP is already shown in Fig. 12 in red colour. Although this spanning cluster has dangling ends, it looks similar to the spanning cluster of Fig. 4 (a) and (b) shown in red. The mass of the spanning cluster of RBMWP at $\omega=\omega_{c}$ is supposed to scale with the lattice size $L$ as $M_{b,\rm max}(\omega_{c},L)\sim L^{d^{\prime}_{f}}$. In Fig. 16, we plot the average mass $M_{b,\rm max}$ of the spanning cluster against $L$ in a double logarithmic scale for $p=0.44,0.47\text{ and }0.50$. The black straight line is the line of linear regression with slope $d^{\prime}_{f}=1.895\pm 0.004$. It is observed that the average mass of the spanning cluster $M_{b,\rm max}$ at $\omega=\omega_{c}$ are very close to each other for all values of $p$. They become indistinguishable on a log scale. Though $\omega_{c}$s are different for different $p$, as the system becomes critical for each $p$, not only does the fractal dimension become the same as that of ordinary site percolation, but also the mass of the spanning cluster becomes the same for all $p$. Thus, the fractal dimension $d^{\prime}_{f}=d_{f}$ remains unchanged.
Since the exponents $\nu^{\prime}$, $\tau^{\prime}$ and the fractal dimension $d^{\prime}_{f}$ are similar to those of the ordinary site percolation, RBMWP seems to belong to the same universality class of ordinary site percolation as it is observed for mixed wet percolation.
@@ 4.6 Finite size scaling
Now, we test the finite size scaling of the geometrical quantities. First, we consider the order parameter $B_{\infty}(p,L)=L^{-\beta/\nu}\widetilde{B}_{\infty}\left[(p-p_{c})L^{1/\nu}\right]$ as given in Eq. 14. In Fig. 17 (a), we plot the scaled order parameter $B_{\infty}(p,L)L^{\beta/\nu}$ against the scaled variable $(p-p_{c})L^{1/\nu}$ taking $\beta/\nu=5/48$ and $41/\nu=3/$ 4 as that of ordinary site percolation . The variations of $B_{\infty}$ against $p$ for four different system sizes $L=256$, $512$, $1024$ and $2048$ are shown in the inset. A good collapse of data of all four system sizes onto a single curve is found for the scaled order parameter $B_{\infty}(p,L)L^{\beta/\nu}$ against the scaled variable $(p-p_{c})L^{1/\nu}$. Thus, the finite size scaling form assumed for $B_{\infty}(p,L)$ is found to be valid with the right values of the critical exponents $\beta$ and $\nu$.
Refer to caption Refer to caption
(a) (b)
@@
Figure 17: (a) Plot of scaled order parameter $B_{\infty}(p,L)L^{\beta/\nu}$ against the scaled variable $(p-p_{c})L^{1/\nu}$ taking $\beta/\nu=5/48$ and $41/\nu=3/$ 4. Plots of $P_{\infty}(p,L)$ against $p$ for different system sizes $L$ are in the inset. A good collapse of data is observed with the chosen values of the critical exponents. (b) Plot of scaled fluctuation $\chi(p,L)L^{-\gamma/\nu}$ against the scaled variable $(p-p_{c})L^{1/\nu}$ taking $\gamma/\nu=43/24$ and $41/\nu=3/$ 4. In the inset, plots of $\chi(p,L)$ against $p$ for different system sizes $L$ are shown. A good collapse of data is observed, which indicates the right choice of the critical exponent’s values.
@@
Next, we study the FSS of the fluctuation $\chi$ in the order parameter. The scaling form of the fluctuation is $\chi(p,L)=L^{\gamma/\nu}\widetilde{\chi}\left[(p-p_{c})L^{1/\nu}\right]$ as given in Eq. 15. In Fig. 17 (b), we plot the scaled fluctuation $\chi(p,L)L^{\gamma/\nu}$ against the scaled variable $(p-p_{c})L^{1/\nu}$ taking $\gamma/\nu=43/24$ and $41/\nu=3/$ 4. The plots of $\chi(p,L)$ against $p$ for four different system sizes $L=256$, $512$, $1024$ and $2048$ are shown in the inset. It can be seen that the fluctuation around $p_{c}$ is diverging as $L$ is increasing. The scaled fluctuations against the scaled variable are found to have a good collapse of data onto a single curve for all four system sizes. Thus, the finite size scaling form assumed for $\chi(p,L)$ is found to be valid. Hence, the finite size scaling exponent $\gamma/\nu$ that describes the scaling of $\chi(p,L)$ with $L$ is the same as that of the ordinary percolation. The scaling relation $2\beta/\nu+\gamma/\nu=d$ and all other scaling relations involving $\gamma$ and $\nu$ are automatically valid.
Figure 18: Plot of scaled fluctuation $\chi^{\prime}(\omega,L)L^{-\gamma^{\prime}/\nu^{\prime}}$ against the scaled variable $(\omega-\omega_{c})L^{1/\nu^{\prime}}$ taking $\gamma^{\prime}/\nu^{\prime}=43/24$ and $41/\nu^{\prime}=3/$ 4. Plots of $\chi(\omega,L)$ against $\omega$ for different system sizes $L$ are shown in the inset. With the chosen values of the critical exponents, a good collapse of data is observed for the scaled fluctuation against the scaled variable.
@@
Finally, we verify the FSS form of the fluctuation $\chi^{\prime}$ in the order parameter of RBMWP. The scaling form of the fluctuation is $\chi^{\prime}(\omega,L)=L^{\gamma^{\prime}/\nu^{\prime}}\widetilde{\chi^{\prime}}\left[(\omega-\omega_{c})L^{1/\nu^{\prime}}\right]$. We plot The scaled fluctuation $\chi^{\prime}(\omega,L)L^{\gamma^{\prime}/\nu^{\prime}}$ is plotted against the scaled variable $(\omega-\omega_{c})L^{1/\nu^{\prime}}$ in Fig. 18 taking $\gamma^{\prime}/\nu^{\prime}=43/24$ as that of ordinary site percolation . It is already found that $41/\nu^{\prime}=1/\nu=3/$ 4. The variation of the fluctuation $\chi^{\prime}(\omega,L)$ against $\omega$ for different $L$ ($256$, $512$, $1024$ and $2048$) is shown in the inset. We found a good collapse of data onto a single curve for all four system sizes. Thus, the finite size scaling form assumed for $\chi^{\prime}(\omega,L)$ is found to be valid. Thus, the finite size scaling exponent $\gamma^{\prime}/\nu^{\prime}$ that describes the scaling of $\chi(\omega,L)$ with $L$ is the same as that of the ordinary site percolation and mixed wet percolation. Hence, the scaling relations involving $\gamma$’, $\nu^{\prime}$ and other exponents will be automatically valid.
@@
It is surprising that mixed wet percolation with a completely different cluster geometry than that of bond percolation exhibits critical behaviour with exactly the same values of the critical exponents as well as the fractal dimension. The non-linear scaling given in Eq. 2 ($\phi(\rho)\sim\left|\Delta\rho\right|^{\pm q}$) does not lead to any new critical exponent. Moreover, even after imposing randomness to the bond occupation in the RBMWP, the critical behaviour does not change. Both the models, mixed wet percolation and RBMWP, belong to the same universality class as ordinary percolation. Importantly, the loop percolation in a different way studied by Pfeiffer and Rieger [34] is also found to belong to the ordinary percolation universality class. However, spiral lattice animals [35] and spiral percolation [36] with loops in either clockwise or anticlockwise directions belong to the spiral percolation universality class, different from ordinary percolation.

### v2

@@ 4.4 Fractal dimension
Since the exponents $\tau$ and $\sigma$ of the cluster size distribution are the same as those of the ordinary site percolation, all other exponents related to different moments of the $n_{b}(p)$ would be the same. However, it would be interesting to see whether the fractal dimension will be that of the percolation hull ziff1986test or that of the ordinary site percolation clusters. We present here an independent measurement of the fractal dimension of the spanning perimeter bond clusters. A typical spanning perimeter bond cluster (in red) is shown in Fig.4(a). The spanning cluster looks very different from a spanning ordinary bond percolation cluster in Fig.4(b). However, the cluster has holes of all possible sizes. It is tortuous and looks self-similar. The mass of the spanning perimeter bond clusters at $p=p_{c}$ is supposed to scale with the system size $L$ as $B_{\rm max}(p_{c},L)\sim L^{d_{f}}$. Note that $B_{\rm max}(\rho_{c},L)$ and $B_{\rm max}(p_{c},L)$ are the same, as they correspond to the same critical point. In Fig. 7, the average mass $B_{\rm max}$ of the spanning cluster (shown in black circles) is plotted against $L$ on a log-log scale. The black straight line of slope $d_{f}=91/48$ is a guide to the eye. The data points follow the black straight line quite nicely. The blue squares represent the average mass $S_{\rm max}$ of the spanning clusters of occupied sites on the primal lattice. Those site clusters represent the nnn site percolation spanning clusters. The blue line is parallel to the red line and has a slope $d_{f}=91/48$. Thus, the fractal dimension $d_{f}$ of the mixed wet percolation spanning cluster and that of the nnn site percolation are the same as that of the ordinary (nn) site percolation spanning cluster. It can be seen that $B_{\rm max}(L)>S_{\rm max}(L)$ as $\rho_{c}>p_{c}$. Therefore, even if there is a mass difference at a given $L$, both of them scale with the same fractal dimension. Since $d_{f}$ and other exponents are the same as those of percolation, all the hyper-scaling relations are satisfied for the perimeter bond cluster percolation. It is surprising that the fractal dimension of the perimeter loops of MWP is different from the hull fractal dimension ($d_{h}=7/4$) and identical to that of the bulk ($d_{f}=91/48$). A closer look at the infinite cluster of MWP reveals that it is essentially made of several perimeter loops connected by knots and forms a giant cluster that covers the whole lattice instead of being the perimeter of a single large cluster of occupied (unoccupied) sites. The giant perimeter loop cluster has almost the same tortuosity or randomness as that of the percolation cluster, and that results in a bulk fractal dimension.
@@ 4.5 Random bond mixed wet percolation (RBMWP) model and results
In Fig. 8, we present a typical morphology of RBMWP on $100\times 100$ square dual lattice at $p=0.5$ and $\omega=0.935$. The RBMWP clusters do not look very different from the mixed wet percolation clusters shown in Fig. 4(a) except for some dangling bonds. The clusters in RBMWP need not always be a combination of closed loops.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 9: (a) Plot of $\omega_{c}(L)$ versus $L^{-1/\nu^{\prime}}$ taking $\nu^{\prime}=4/3$, the correlation length exponent of percolation. This is done for values of $p=0.44,0.47$ and $0.50$. (b) Plot of $\omega_{c}$ for various values of $p$ between $p_{c}$ and $1-p_{c}$. (c) Plot of $\Pi^{\prime}_{b}(\omega_{c},L)$ versus $L$ in double logarithmic scale for $p=0.44,0.47$ and $0.50$. The correlation length exponent is found as $1/\nu^{\prime}\approx 0.74\pm 0.004$ for all three cases.
@@
$\Pi_{b}(\omega,L)=\widetilde{\Pi}_{b}\left[(\omega-\omega_{c})L^{1/\nu^{\prime}}\right]\;,$ (24)
@@
$\Pi_{b}(\omega_{c}(L),L)=\widetilde{\Pi}_{b}\left[(\omega_{c}(L)-\omega_{c})L^{1/\nu^{\prime}}\right]=1/2\;,$ (25)
@@
$\omega_{c}(L)=\omega_{c}+AL^{-1/\nu^{\prime}}\;,$ (26)
@@
$\Pi^{\prime}_{b}(\omega_{c},L)=aL^{1/\nu^{\prime}}\;,$ (27)
where $a=\widetilde{\Pi}^{\prime}_{b}[0]$ and the prime () indicates the derivative with respect to $\omega$. The derivatives are measured at the critical thresholds $\omega_{c}$ by fitting a third-degree polynomial around $\omega_{c}$ for each $L$. In Fig. 9(c), $\Pi^{\prime}_{b}$ is plotted against $L$ in a double logarithmic scale. The slope that indicates $1/\nu^{\prime}$ was determined by linear regression and was found to be $1/\nu^{\prime}=0.740\pm 0.004$. That is very close to that of the ordinary percolation one, i.e.; $\nu^{\prime}=\nu=4/3$.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 10: (a) Plot of cluster size distribution $n_{b}(\omega_{c})$ against the size of the clusters $b$ at the threshold $\omega_{c}=0.9350$ for $p=0.5$ on a double logarithmic scale for a lattice size $L=1000$. The red line of slope $91/48$ is a guide to the eye. (b) Plot of the scaled cluster size distribution $n_{b}(\omega)/n_{b}(\omega_{c})$ against the scaled variable $(\omega-\omega_{c})b^{\sigma}$ for $p=0.5$ on a lattice of size $L=2000$. A good collapse of data is obtained for $\sigma^{\prime}=\sigma=36/91$. (c) Plot of mass of the spanning cluster $M_{b,\rm max}$ against lattice size $L$ on a double logarithmic scale. The black straight line is the line of linear regression with slope $d_{f}=1.895\pm 0.004$.
@@
$n_{b}(\omega)=b^{-\tau^{\prime}}\mathcal{N}_{b}\left[(\omega-\omega_{c})b^{\sigma^{\prime}}\right]\;,$ (28)
@@
$n_{b}(\omega_{c})\sim b^{-\tau^{\prime}}\;,$ (29)
where $b$ is the size of the clusters and $\tau^{\prime}$ is the cluster size distribution exponent. In Fig. 10 (a), $n_{b}(\omega_{c})$, for $p=0.5$ and $\omega_{c}=0.9350$, is plotted against the cluster size $b$ on a double logarithmic scale. The exponent $\tau^{\prime}$ measured by linear regression of the data points is found to be $\tau^{\prime}=2.058\pm 0.005$. It is very close to the exponent $\tau=187/91$ as that of ordinary site percolation, shown by a red line as a guide to the eye. It seems $\tau^{\prime}$ of RBMWP is the same as that $\tau$ of ordinary site percolation, as well as mixed wet percolation.
@@
$\frac{n_{b}(\omega)}{n_{b}(\omega_{c})}\sim\mathcal{N}_{b}\left[(\omega-\omega_{c})b^{\sigma^{\prime}}\right]\;,$ (30)
@@
Although $\nu^{\prime}$ and $\tau^{\prime}$ were very close to corresponding exponents of ordinary site percolation, we will verify the fractal dimension of the spanning clusters of RBMWP. A typical spanning cluster of RBMWP is already shown in Fig. 8 in red colour. Although this spanning cluster has dangling ends, it looks similar to the spanning cluster of Fig. 4 (a) shown in red. The mass of the spanning cluster of RBMWP at $\omega=\omega_{c}$ is supposed to scale with the lattice size $L$ as $M_{b,\rm max}(\omega_{c},L)\sim L^{d^{\prime}_{f}}$. In Fig. 10(c), we plot the average mass $M_{b,\rm max}$ of the spanning cluster against $L$ in a double logarithmic scale for $p=0.44,0.47\text{ and }0.50$. The black straight line is the line of linear regression with slope $d^{\prime}_{f}=1.895\pm 0.004$. It is observed that the average mass of the spanning cluster $M_{b,\rm max}$ at $\omega=\omega_{c}$ is very close to each other for all values of $p$. They become indistinguishable on a log scale. Though $\omega_{c}$s are different for different $p$, as the system becomes critical for each $p$, not only does the fractal dimension become the same as that of ordinary site percolation, but also the mass of the spanning cluster becomes the same for all $p$. Thus, the fractal dimension $d^{\prime}_{f}=d_{f}$ remains unchanged.
Since the exponents $\nu^{\prime}$, $\tau^{\prime}$ and the fractal dimension $d^{\prime}_{f}$ are similar to those of the ordinary site percolation, RBMWP seems to belong to the same universality class of ordinary site percolation as it is observed for mixed wet percolation. Though these models have different cluster size distributions $n_{b}(p)$ and critical masses $B_{\mathrm{max}}(p_{c},L)$ of the infinite clusters, their scaling behaviour remains the same with their respective cluster size $b$ or with the system size $L$. That is why all models are in the same universality class.
@@ 4.6 Finite size scaling
Now, we test the finite-size scaling of the geometrical quantities. First, we consider the order parameter $B_{\infty}(p,L)=L^{-\beta/\nu}\widetilde{B}_{\infty}\left[(p-p_{c})L^{1/\nu}\right]$ as given in Eq. 14. In Fig. 11 (a), we plot the scaled order parameter $B_{\infty}(p,L)L^{\beta/\nu}$ against the scaled variable $(p-p_{c})L^{1/\nu}$ taking $\beta/\nu=5/48$ and $1/\nu=3/4$ as that of ordinary site percolation . The variations of $B_{\infty}$ against $p$ for four different system sizes $L=256$, $512$, $1024$ and $2048$ are shown in the inset. A good collapse of data of all four system sizes onto a single curve is found for the scaled order parameter $B_{\infty}(p,L)L^{\beta/\nu}$ against the scaled variable $(p-p_{c})L^{1/\nu}$. Thus, the finite size scaling form assumed for $B_{\infty}(p,L)$ is found to be valid with the right values of the critical exponents $\beta$ and $\nu$.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
@@
Figure 11: (a) Plot of scaled order parameter $B_{\infty}(p,L)L^{\beta/\nu}$ against the scaled variable $(p-p_{c})L^{1/\nu}$ taking $\beta/\nu=5/48$ and $1/\nu=3/4$. Plots of $B_{\infty}(p,L)$ against $p$ for different system sizes $L$ are in the inset. A good collapse of data is observed with the chosen values of the critical exponents. (b) Plot of scaled fluctuation $\chi(p,L)L^{-\gamma/\nu}$ against the scaled variable $(p-p_{c})L^{1/\nu}$ taking $\gamma/\nu=43/24$ and $1/\nu=3/4$. In the inset, plots of $\chi(p,L)$ against $p$ for different system sizes $L$ are shown. A good collapse of data is observed, which indicates the right choice of the critical exponent’s values. (c) Plot of scaled fluctuation $\chi^{\prime}(\omega,L)L^{-\gamma^{\prime}/\nu^{\prime}}$ against the scaled variable $(\omega-\omega_{c})L^{1/\nu^{\prime}}$ taking $\gamma^{\prime}/\nu^{\prime}=43/24$ and $1/\nu^{\prime}=3/4$. Plots of $\chi(\omega,L)$ against $\omega$ for different system sizes $L$ are shown in the inset. With the chosen values of the critical exponents, a good collapse of data is observed for the scaled fluctuation against the scaled variable.
@@
Next, we study the FSS of the fluctuation $\chi$ in the order parameter. The scaling form of the fluctuation is $\chi(p,L)=L^{\gamma/\nu}\widetilde{\chi}\left[(p-p_{c})L^{1/\nu}\right]$ as given in Eq. 15. In Fig. 11 (b), we plot the scaled fluctuation $\chi(p,L)L^{\gamma/\nu}$ against the scaled variable $(p-p_{c})L^{1/\nu}$ taking $\gamma/\nu=43/24$ and $1/\nu=3/4$. The plots of $\chi(p,L)$ against $p$ for four different system sizes $L=256$, $512$, $1024$ and $2048$ are shown in the inset. It can be seen that the fluctuation around $p_{c}$ is diverging as $L$ increases. The scaled fluctuations against the scaled variable are found to have a good collapse of data onto a single curve for all four system sizes. Thus, the finite-size scaling form assumed for $\chi(p,L)$ is found to be valid. Hence, the finite size scaling exponent $\gamma/\nu$ that describes the scaling of $\chi(p,L)$ with $L$ is the same as that of the ordinary percolation. The scaling relation $2\beta/\nu+\gamma/\nu=d$ and all other scaling relations involving $\gamma$ and $\nu$ are automatically valid.
@@
Finally, we verify the FSS form of the fluctuation $\chi^{\prime}$ in the order parameter of RBMWP. The scaling form of the fluctuation is $\chi^{\prime}(\omega,L)=L^{\gamma^{\prime}/\nu^{\prime}}\widetilde{\chi^{\prime}}\left[(\omega-\omega_{c})L^{1/\nu^{\prime}}\right]$. We plot The scaled fluctuation $\chi^{\prime}(\omega,L)L^{\gamma^{\prime}/\nu^{\prime}}$ is plotted against the scaled variable $(\omega-\omega_{c})L^{1/\nu^{\prime}}$ in Fig. 11(c) taking $\gamma^{\prime}/\nu^{\prime}=43/24$ as that of ordinary site percolation . It is already found that $1/\nu^{\prime}=1/\nu=3/4$. The variation of the fluctuation $\chi^{\prime}(\omega,L)$ against $\omega$ for different $L$ ($256$, $512$, $1024$ and $2048$) is shown in the inset. We found a good collapse of data onto a single curve for all four system sizes. Thus, the finite-size scaling form assumed for $\chi^{\prime}(\omega,L)$ is found to be valid. Thus, the finite size scaling exponent $\gamma^{\prime}/\nu^{\prime}$ that describes the scaling of $\chi(\omega,L)$ with $L$ is the same as that of the ordinary site percolation and mixed wet percolation. Hence, the scaling relations involving $\gamma$’, $\nu^{\prime}$ and other exponents will be automatically valid.
@@
It is surprising that mixed wet percolation with a completely different cluster geometry than that of bond percolation exhibits critical behaviour with exactly the same values of the critical exponents as well as the fractal dimension. The non-linear scaling given in Eq. 2 ($\phi(\rho)\sim\left|\Delta\rho\right|^{\pm q}$) does not lead to any new critical exponent. Moreover, even after imposing randomness to the bond occupation in the RBMWP, the critical behaviour does not change. Both the models, mixed wet percolation and RBMWP, belong to the same universality class as that of ordinary percolation. Interestingly, the loop percolation in a different way studied by Pfeiffer and Rieger pfeiffer2003critical is also found to belong to the ordinary percolation universality class. However, spiral lattice animals santra1989statistics and spiral percolation santrs1992spiral with loops in either clockwise or anticlockwise directions belong to the spiral percolation universality class, different from ordinary percolation.